# Low-Overhead Channel Estimation via 3D Extrapolation for TDD mmWave Massive MIMO Systems Under High-Mobility Scenarios

### v1

@@ VI Conclusion
This paper proposed a spatial, frequency, and temporal channel extrapolation framework for TDD mmWave massive MIMO-OFDM systems to systematically reduce the pilot training overhead under high-mobility scenarios. Specifically, two neural networks, namely the KDD-SFCEN and the TUDCEN, were proposed to reduce the spatial-frequency domain pilot training overhead and to reduce the temporal domain pilot training overhead, respectively. Extensive numerical results demonstrated that the proposed spatial, frequency, and temporal channel extrapolation framework can effectively reduce the spatial-frequency domain pilot training overhead by more than $4$ times via spatial-frequency channel extrapolation and further reduce the temporal domain pilot training overhead by additional $4$ times and significantly improve the system’s spectral efficiency via enlarging the pilot signal period with slot-level channel extrapolation.

### v2

@@ VI Conclusion
This paper proposed a spatial, frequency, and temporal channel extrapolation framework for TDD mmWave massive MIMO-OFDM systems to systematically reduce the pilot training overhead under high-mobility scenarios. Specifically, two neural networks, namely the KDD-SFCEN and the TUDCEN, were proposed to reduce the spatial-frequency domain pilot training overhead and to reduce the temporal domain pilot training overhead, respectively. Extensive numerical results demonstrated that the proposed 3D channel extrapolation framework can effectively reduce the spatial-frequency domain pilot training overhead by $4$ times via spatial-frequency channel extrapolation and further reduce the temporal domain pilot training overhead by additional $4$ times via reducing the times of single-slot channel estimations with slot-level channel extrapolation.
VII Acknowledgement
This work was performed in part at SICC which is supported by SKL-IOTSC, University of Macau.